The average speed of nitrogen gas molecular at 25 degree celsius is 515 m/s. (a) what is the average speed of a He molecules at 25 degree celsius. (b) a CO2 gas atom at 25 degree celsius.
1
Expert's answer
2020-11-26T10:54:05-0500

ACCORDING TO :

average speed of nitrogen gas molecular at 25 degree celsius is 515 m/s. 

we know the Vrms of N2 is 515 m/s.


Now we calculate the Vrms for He2 and CO2

Now we calculate Vrms of He2

Vrms = "\sqrt\frac{3RT}{M}" R= 8.314J⋅K−1⋅mol−1.

and T = 273 + 25⁰C = 298K

Mass of He2 = 8.0052 × "10^-\u00b3"

Vrms = "\sqrt\frac{3×8.314×298}{8.0052×(10)^-³}"

Vrms = 963.57 m/s

This is the average speed of He molecule is 963.57 m/s.


Now we calculate Vrms ofCO2

Vrms ="\sqrt\frac{3RT}{M}" R =8.314J⋅K−1⋅mol−1.

and T = 273 + 25⁰C = 298K

Mass of CO2 = 44.01 × "10^-\u00b3"

Vrms = "\sqrt\frac{3×8.314×298}{44.01×10^-³}"

Vrms = 410.95 m/s

This is the average speed of CO2 atom is 410.95 m/s.